Fermi National Laboratory
 

Accelerator Update for January 19, 2009 - January 21, 2009

 

Monday Jan 19

The day shift began with the Tevatron (TeV) in access, with Operations monitoring the Recycler stash (83.9E10), and the Antiproton Source stacking (28.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments in access.

At 8:32 AM, a Booster RF station (BRF8) tripped off. It reset.

CDF came out of access at 8:46 AM.

D0 came out of access at 8:55 AM.

At 9:19 AM, the Booster East Anode Power Supply tripped off. Operators discovered that BRF5 was the cause of the trip. They bypassed the station and reset the supply, but with two RF stations off, Operations had to lower the Booster beam intensity.

At 10:50 AM, LRF5 tripped off for the first of two times in three minutes. It reset.

Tevatron Electron Lens (TEL1) experts reported at 11:46 AM that they were still investigating.

Operations reported at 12:03 PM that the quench previously caused by TEL1 was a controls clock problem.

An RF expert brought BRF5 back online at 12:25 PM.

At 12:35 PM, BRF3 and 4 tripped off. They reset.

TEL1 was brought online at 12:53 PM.

At 1:27 PM, the TeV suffered a sector F1 quench. There was no beam in the machine.

Operations reported at 3:55 PM that the Recycler had stashed antiprotons seven times during the day shift.

At 4:45 PM, Operations reported that they found the TEL1 clock set wrong and this was the cause of the quench.

Operations put the TeV through a dry squeeze at 5:25 PM.

At 5:31 PM, a Linac quadrupole power supply (QPS111) tripped off. It reset.

Operations put the TeV into shot setup at 6:37 PM.

At 6:48 PM, the TEL1 began acting oddly. Operations put the shot setup on hold while an expert investigated.

At 7:19 PM, Operations halted the shot setup and put the TeV into a dry squeeze.

At 8:36 PM, the Linac Klystron RF stations 1-7 all tripped off on a LCW flow indication. Linac water skid #3 had tripped off. An operator tapped on the water skid flow meter and the alarm cleared.

At 9:58 PM, a MI kicker (KPS5) tripped off. It reset.

At 10:14 PM, KRF5 tripped off. It reset.

At 10:22 PM, Operations reported that a TEL expert was coming in to switch the abort cleaning duties to TEL2.

Operations reported at 11:58 PM that the Recycler had stashed six times during the evening shift.

Tuesday Jan 20

The midnight shift began with experts working on reconfiguring TEL2 to clean the abort gap, with Operations monitoring the stash (375.5E10), with Pbar stacking (30.0E10), and with NuMI and MiniBooNE taking beam.

Operations put the TeV through a dry squeeze around midnight.

TEL experts reported that they had TEL2 ready to go. Operations put the TeV through a wet squeeze at 4 AM.

Operations made changes to the TeV orbit at 5:30 AM and the put the TeV through another wet squeeze.

Operations put the TeV into shot setup at 6:39 AM.

Operations reported at 8:23 AM that the Recycler had stashed twice during the midnight shift.

Operations established store 6745 at 8:24 AM with an initial luminosity of 352.04E30.

At 8:41 AM, Operations halted beam to MiniBooNE due to a sump pump problem in the MI-12B tunnel.

At 8:44 AM, Operations reported two empty transfers of antiprotons due to an Accumulator RF (ARF4) station not ramping properly. Operators investigated and found that the Flux capacitor was not responding properly.

Operations prepared MI-12B area for access at 10:10 AM.

Operations reported at 10:59 AM that the Pbar Oscillation Overthruster program (Pbar Autotune) was not communicating with the P1 transfer beam line BPMs. Operators rebooted the IBPMP1.

A MI expert began a study around 11:15 AM.

At 12:06 PM, a MI kicker (KPS5S) tripped off. It reset.

A MI expert began a 30-minute MI-8 study at 2:20 PM.

TeV experts reported at 3:07 PM that they had made lattice adjustments around D0.

Operations reported at 3:57 PM that the Recycler had stashed eight times during the day shift.

At 4:48 PM, a MI kicker (KPS62A) tripped off. It reset.

At 5:30 PM, a Booster vacuum pump (B:IPR24) tripped off and caused BRF16 and 17 to trip. Everything reset.

At 5:51 PM, B:IPR24 tripped off and caused BRF13 and 14 to trip off. Everything reset.

At 7:46 PM Operations reported that the Recycler VSA had failed for the second time and would not recover. Operations contacted an expert who suggested operators turn the VSA off for 15 minutes and then turn it back on. This worked.

At 8:42 PM, a MI kicker (KPS5S) tripped off. It reset.

Operations terminated store 6745 at 10:01 PM.

Operations put the TeV into shot setup at 10:11 PM.

Operations established store 6746 at 11:38 PM with an initial luminosity of 337.95E30.

At 11:46 PM, a MI kicker (KPS5S) failed to fire.

At 11:51 PM, Operations reported that TEL2 did not come on. Operations contacted an expert.

Operations reported at midnight that the Recycler had stashed seven times during the evening shift.

Wednesday Jan 21

The midnight shift began with Operations monitoring store 6746 (277.64E30) and the stash (74.2E10), with Pbar stacking (17.3E10), and with NuMI and MiniBooNE taking beam.

At 3:07 AM, a MI power supply’s (I:V714) current drifted low and caused the P1 transfer line beam permit to trip off. Everything reset.

At 5:12 AM, a MI kicker (KPS62A) tripped off. It reset.

At 5:13 AM, a MI sextupole power supply (I:SEXVPS) tripped off. It reset.

At 5:50 AM, KRF3 tripped off. It reset.

At 7:13 AM, KRF1 tripped off for the first of two times in 20 minutes. It reset.

At 7:46 AM, a MI kicker (KPS5S) failed to fire.

Operations reported at 7:56 AM that the Recycler had stashed eight times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/21/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory